OpenStudy (anonymous):

Medals!! Tabitha is trying to find the equation of a line perpendicular to y = 1 over 2x − 5 in slope-intercept form that passes through the point (2, −7). Which of the following equations will she use?

OpenStudy (anonymous):

a) y − (−7) = 1 over 2(x − 2) b) y − 2 = 1 over 2(x − (−7)) c) y − (−7) = −2(x − 2) d) y − 2 = −2(x − (−7))

OpenStudy (anonymous):

First find the slope of the line. You are given the slope of the perpendicular line.

OpenStudy (anonymous):

To get the slope of the line, you take the negative reciprocal of the perpendicular slope.